b) Explain possible actions that might be taken to reduce the length of that cycle, 

and the possible disadvantages of each.

c) Assume that the company is negotiating to do business with a new supplier 

who has offered credit terms of 3/15, net 30. The financial manager is planning 

to delay payment for an additional 10 days i.e. to only settle the account after 

40 days. The current bank overdraft rate for the firm is 25% per annum.

Calculate the effective cost of finance provided by this supplier and comment 

on the financial manager’s plans. [3 Marks]

d) An aspect of working capital policy which requires managerial attention is the 

manner in which the items are financed. Discuss aggressive policy in this 

regard.


1
Expert's answer
2021-06-14T15:06:39-0400

b) Adequate cash flow is extremely essential to the success of the business. It is a vital metric in a company’s well-being. The quantity of accessible cash of the business has at hand at any particular time impacts both business’s day-to-day and long-term processes and operations. Poor cash flow makes the business’s daily commercial processes and operations more challenging and it also harmfully impacts the time-frame on paying creditors. Decreasing the cycle length can even enhance the company’s bottom line through interest savings.

c)


= 3% / (100% - 3%) x 360 / (30 - 15) = 74.23%

The active budget of investment delivered by this contractor = 74.23%.

Executive's strategy: He or she will delay the disbursement by further 10 days. In that circumstance the closing due date becomes = 30 + 10 = 40 days.

Charge of not taking discount then becomes = = 3% / (100% - 3%) x 360 / (40 - 15) = 44.54% > 25% = the bank's overdraft rate. Therefore, the executive's strategy is not ideal, he or she must borrow and pay the contractor on 15th day and revel in the discount.

d) Aggressive policy: Temporary coffers are employed to fund durable assets. Illustration will be: Utility of temporary credits / overdraft// working capital loans to fund the durable fixed resources such as machinery and equipment. This brings about resource liability maturity incongruity.

Moderate: Temporary capitals are employed to fund short-term assets, while long-terms capitals are employed to fund the durable resources. This is extensively monitored working capital management strategy. This brings about usage of working capital finances / rotating capacity / overdraft towards funding / bankrolling receivables and or register.

Conservative: Utility of long-term coffers towards temporary resources, the firm borrows on continuing foundation to fund the working capital necessities, which is sub-optimal.